OpenStudy (anonymous):

Evaluate the expression then convert your answer to decimal form : (2/3)^2=____ ?? help please

OpenStudy (anonymous):

.4 repeating

OpenStudy (anonymous):

(2/3)^2 = (2/3)(2/3) = 4/9 now its just a matter of dividing, and you get 0.4444.. like josh said